A boat travels 42 miles up the river in the same amount of time it takes to travel 58 miles down the same river. If the current is 4 miles per hour, what is the speed of the boat in still water?
Answer by TimothyLamb(4379)

up:
s = d/t
b - c = d/t
b - c = 42/t
---
down:
s = d/t
b + c = d/t
b + c = 58/t
---
c = 4
---
b - 4 = 42/t
b + 4 = 58/t
---
b = 42/t + 4
b = 58/t - 4
---
42/t + 4 = 58/t - 4
8 = 58/t - 42/t
8 = (58 - 42)/t
8 = 16/t
t = 16/8
t = 2 hours
---
b + 4 = 58/2
b = 58/2 - 4
b = 29 - 4
---
answer:
b = 25 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php